### v5

@@ 4 Results
All models were fitted with a training loop comprised of 100 epochs with a batch size of 128. All experiments adopted the recommended learning rate for the Adam optimizer of 0.001 [50].
@@ 5 Conclusion
We introduced in this work a newly designed Network in Network block that substitutes the attention block used traditionally in transformer architectures by extending the token mixing approach presented in the MLP-Mixer to function as a gating signal generator and taking advantage of the gating mechanism to introduce dynamic behavior in the newly proposed block, thus enhancing the static weight approach of the MLP-Mixer by utilizing its layers as a sub-unit network within the outer network formulation. The experimental results show that in comparison to baseline architectures that were chosen from different families of transformer formulations, our proposed block significantly outperforms, showing noticeable improvements on those baselines, specifically showing a great enhancement of accuracy compared to the standalone MLP-Mixer architecture that acts as a sub-unit, which validates that our proposal, with its dynamic gating of the upstream representation, properly enhances and circumvents the shortcoming of the static weight approach of the standalone MLP-Mixer while still providing more simplicity of operations in contrast to the vanilla ViT transformer architecture. Future directions of this work are to investigate a multitude of sub-unit network selections, aiming for further enhancements and capabilities.

### v6

@@ 4 Results
All models were fitted with a training loop comprised of 100 epochs with a batch size of 128. All experiments adopted the recommended learning rate for the Adam optimizer of 0.001 [50], other hyper-parameters such as patch size and token dimension were chosen so that it saturates the hardware capacity provided by the available computer system.
@@
The performance of deep learning models is highly dependent on the low-level hardware details and software optimizations [51]; the timing of execution shows significant sensitivity to the interactions between micro-architectural and execution characteristics such as caches and RAM configurations.We have performed per-sample inference time measurements on the selected baseline architectures and the proposed architecture of this work conducted as relative performance measures in relation to the hardware system available for the purposes of this work. Table 3 illustrates the obtained inference time results after performing the experimentation on MNIST, CIFAR-10 and CIFAR-100 datasets applied to the baseline architectures and NiNformer architecture.
Table 3: Experimental per-sample inference time measured in nano-seconds obtained on the utilized dataset.
Models Data sets
MNIST CIFAR-10 CIFAR-100
ViT 141.62 142.64 141.07
MlpMixer 132.68 103.53 104.24
LocalViT 139.65 127.76 115.32
NiNformer (ours) 135.00 104.64 105.37
The obtained measurements are in support of the design goals of our proposals, as the inference time of our work adds a low inference time cost on the MLP-Mixer, which is used as a subunit within our work.Taking the measurements on CIFAR-10 as a reference, the execution time cost is only an additional 1%, while the improvement in accuracy over the MLP-Mixer is significant at 16%. Extending the comparison to the other baselines, our proposal shows a wide gain in accuracy of 24% and 16% for ViT and Local-ViT respectively, while also gaining a significant improvement margin in inference time measurements of 36% and 22% for ViT and Local-ViT, respectively.The results are in high accordance with the hypothesis introduced in this work of formulating a novel computational block that enhances the capacity and capability of linear alternatives to the attention mechanism while maintaining the properties of efficiency and fast execution margins over the traditional formulation of Transformers.
@@ 5 Conclusion
This work introduced a newly designed Network in Network block that substitutes the attention block traditionally utilized in designing transformer architectures. The proposed efficient and highly performing block extends the token mixing approach presented in the MLP-Mixer to function as a gating signal generator and takes advantage of the gating mechanism to introduce dynamic token processing. The new mechanism of our proposal presents an enhancement of the static weight approach of the MLP-Mixer by utilizing its layers as a sub-unit network incorporated within a gating function of an outer network formulation. The experimental results show that our proposed block significantly outperforms the baseline architectures, offering noticeable improvements on the selected baselines, specifically showing a great enhancement of accuracy compared to the standalone MLP-Mixer architecture that acts as a sub-unit, validating the assumptions of the proposal introduced in this work positing that a two-level Network in Network organization of the main computational block and employing a dynamic gating of the upstream representation results in a significant enhancement and circumvents the shortcoming of the static weight approach of the standalone MLP-Mixer while still providing more simplicity of operations in contrast to the vanilla ViT transformer architecture. Future directions of this work are to investigate a multitude of sub-unit network selections, aiming for further enhancements and capabilities.